# Heralded Optical Entanglement Generation via the Graph Picture of Linear Quantum Networks

### v2

@@ Abstract
Non-destructive heralded entanglement with photons is a valuable resource for quantum information processing. However, they generally entail ancillary particles and modes that amplify the circuit intricacy. To address this challenge, a recent work (arXiv:2211.04042 (2022)) introduced a graph approach for creating multipartite entanglements with boson subtractions. Nonetheless, it remains an essential intermediate step toward practical heralded schemes: the proposition of heralded subtraction operators in bosonic linear quantum networks. This research establishes comprehensive translation rules from subtraction operators to linear optical operators, which provides a seamless path to design heralded schemes with single photons. Our method begets enhanced or previously unreported schemes for the $N$-partite GHZ state with $2N$ photons, $N$-partite W state with $12N+$ 1 photons and superposition of $N=3$ GHZ and W states with 9 photons. Our streamlined approach can straightforwardly design heralded schemes for multipartite entangled states by assembling the operators according to the guidence of sculpting bigraphs, hence significantly simplifies the quantum circuit design process.
I Introduction
@@
In contrast, heralded schemes [11, 12, 13, 14, 15] adopts a different strategy by incorporating ancillary photons and modes as “heralds” to signal the successful generation of the anticipated target states. This allows for the discrimination of experimental runs that produce the desired target states without directly measuring those states. As a result, the entanglement generated through heralded schemes can be harnessed as a useful resource for quantum computations. However, the heralded schemes generally demand additional particles and modes for the resource of heralding, making them more intricate to design than postselected ones.
To overcome this difficulty of heralded schemes, a graph approach to heralded entanglement generation is proposed in Ref. [16], providing a systematic strategy to design the schemes by boson subtractions. Ref. [16] imposed correspondence relations between elements of bipartite graphs (bigraphs) and many-particle systems including boson annihilation operators, which is an extension of the graph picture introduced in Ref. [8]. To distinguish from other graphical approaches in quantum information and computation [17, 18, 19, 5], we call our approach the linear quantum graph (LQG) picture.
Based on the LQG picture, Ref. [16] provided various schemes to generate the multipartite entanglement of bosons with the sculpting protocol [20, 21], which generates an $N$-partite entangled state by applying tailored $N+K$ single-boson subtraction operators (which is dubbed the sculpting operator) with $K$ ancillary bosons. In the LQG picture, all the sculpting operators are represented as bigraphs (which are dubbed scupting bigraphs). By leveraging the advantageous mathematical properties of the graphs, Ref. [16] has found essential entanglement generation schemes including qubit $N$-partite GHZ and W states, $N=3$ Type 5 entangled states (the superposition of $N=3$ GHZ and W states [22]), and qudit $N$-partite GHZ states.
@@
This work is organanized as follows: Section II explains the concept of sculpting protocol and its linear quantum graph (LQG) picture. Section III provides the translation rules from sculpting bigraphs to linear optical elements with the introduction of heralded subtraction operators (subtractors) in linear optics, by which we can construct heralded linear optical circuits that create GHZ, W, and the superposition of $N=3$ GHZ and W states in Section IV. Section V summarized the physical implications of our research and further applications.
II Sculpting protocol and its linear quantum graph (LQG) picture
@@ II.1 Sculpting protocol
    $\quad(k^{(l)}_{j,s}\in\mathbb{C}~{}\textrm{and}~{}\sum_{j,s}|k^{(l)}_{j,s}|^{2}=1).$ (3)
@@
II.2 Directed graph representation of bosonic systems in LQG picture
In Ref. [16], an undirected bigraph mapping was introduced (Table I of Ref. [16]), which was useful for finding several sculpting schemes in the work. One can consult Appendix A of Ref. [8] for a simple glossary to the graph theory. While the undirected bigraph representation includes all the crucial information on the sculpting operators that can generate entanglement, we need a more comprehensive $directed$ bigraph representation to embrace the initial state and the sculpting operator (see Appendix A of Ref. [16]) in the LQG picture. We explain all our schemes based on the directed bigraph representation in this paper as it has a more direct and intuitive relation with heralded linear optical circuits. In the directed bigraph picture, the complete information on the sculpting protocol from the preparation to the final state is encoded in a graph, which we call a sculpting directed bigraph.
@@ II.2 Directed graph representation of bosonic systems in LQG picture
The first identity of Eq. (II.2) is expressed with directed bigraphs as
@@
[…]
@@
II.3 EPM directed bigraphs for generating entanglement
@@
A crucial property of an EPM bigraph is that the final state by the corresponding sculpting operator is fully determined by the perfect matchings (PMs) of the bigraph (a set of disjoint edges that cover all the vertices in a given bigraph, see Appendix A of Ref. [8]), which we can show using identities (8) and (10) (see Sec. IV of Ref. [16] for the detailed explanation). All the sculpting bigraphs for generating genuine qubit entanglement in Ref. [16] correspond to EPM bigraphs.
@@
Comparing the EPM directed bigraphs in Fig. 1 with the EPM (undirected) bigraphs presented in (20), (27), and (32) of Ref. [16], we can easily see that the RHS of the directed ones in Fig. 1 (a), (b), and (c) without direction are equal to (20), (27), and (32) in Ref. [16]. This is by the fact that the information on the sculpting operators is in the RHS side of the directed graphs, which is solely represented in the undirected graphs in Ref. [16]. Therefore, the essential part in a EPM directed bigraph is the RHS of the graph from the circles as it represents the sculpting operator that generates entanglement.
III Translation rules from sculpting bigraph elements to linear optical elements
Ref. [16] exploited the special property of EPM bigraphs (Property 2 of Section IV in Ref. [16]) to find find schemes for genuinely entangled states, which are enumerated in Fig. 1. To design counterparts in linear optical circuits to the EPM bigraphs that generate the same entagled states, we need to know how each element of the graphs is translated to linear optical operators. A prerequisite for the translation rules from graphs to circuits is to provide heralded linear optical operators (which we name subtractors) that play the role of subtraction operators, because they are the fundamental components of EPM sculpting bigraphs.
@@ 2. $K$ circles to which edges are attached as
III.2 Subtraction operators in linear optics with heralding
@@ 2. Half-wave plates with HWP: $\{H,V\}\leftrightarrow\{D,A\}$ and HWP: $\{H,V\}\leftrightarrow\{A,D\}$
3. $n$-partite multiports that Fourier-transform the spatial mode states of photons ($2$-partite ports correspond to beam splitters)
@@
$\{|0\rangle,|1\rangle\}\cong\{|D\rangle,|A\rangle\},~{}~{}~{}\{|+\rangle,|-\rangle\}\cong\{|H\rangle,|V\rangle\}.$ (24)
@@
    ${\hat{a}}^{\dagger}_{H}{\hat{a}}^{\dagger}_{V}\to{\hat{a}}^{\dagger}_{D}{\hat{a}}^{\dagger}_{A}=\frac{1}{2}({\hat{a}}^{\dagger 2}_{H}-{\hat{a}}^{\dagger 2}_{A})$ (25)
@@
    $\frac{1}{2}({\hat{a}}^{\dagger 2}_{1,H}-{\hat{a}}^{\dagger 2}_{2,V})\to{\hat{a}}^{\dagger}_{1,H}$ (27)
    We can see that the measurement of one photon at the detector 12 corresponds to the subtraction of one photon.
@@
where the incoming wires are attached to $11$ and $22$ by the internal states of the annihilation operators. In the above subtractor, the two open wires go back to the spatial modes from which those wires came and be merged by PBSs as in (23) and (29).
@@
With the subtractors as building blocks, we can enumerate all the elements that construct EPM directed bigraphs and present translation rules of them to linear optical elements. All the translation rules from EPM sculpting bigraph elements to linear optical elements that we have explained are enumerated in Fig. 2. The resultant circuits all exhibit interesting advantages over previously known optical schemes.
@@
IV Linear optical circuits
@@
Following the translation rules, we simply obtain a linear optical circuit for the GHZ state as Fig. 3. Comparing the above circuit with the GHZ sculpting bigraph (Fig. 1 (a)), we can see that the photon paths are connected following the structure of the bigraph. The explicit calculation for showing that the above circuit generates the GHZ state is given in Appendix A. The success probability $P_{suc}$ is $\frac{1}{2^{2N-1}}$ with feed-forward.
We can compare our scheme with other heralded schemes. First, our scheme can be directly applied to any $N$-partite case with only $2N$ photons unlike Ref. [24] that suggests schemes for $N=2$ and 3 based on numerical optimizations or Ref. [23] for the $N=3$ GHZ entanglement generation using fusion gates (see Refs. [29, 30] for the experimental results). In addition, contrary to schemes in Refs. [24, 31] that depend on the dual-rail encoding structure, our scheme can be installed with any kind of qubit encoding (polarization, dual-rail, etc.) based on the sculpting bigraph. See Appendix D for the transformation from polarization encoding circuits to dual-rail encoding circuits. More advantages of our scheme from the viewpoint of the communication security and the robustness against lossy channels are analyzed in Ref. [32].
IV.2 $N$-partite W state (Fig. 1, (b))
@@
A linear optical scheme for the W generating EPM bigraph (Fig. 1, (b)) from the translation rules is as in Fig. 4. Considering each dashed box (spatially overlapped subtractors) as a dot, we can see that the wires are connected following the structure of the W sculpting bigraph. Appendix B provides an explicit calculation for showing that the above circuit generates the W state with $P_{suc}=\frac{1}{2^{2N}}$ with feed-forward.
To our knowledge, no optical heralded W-state scheme for arbitrary $N$ party with single-photon source has been proposed. Since our scheme requires $12N+$ 1 single-photons, it is much more feasible than those schemes which exploit fusion gates and small size of W states as resources for generating $N$-partite W states [25, 26].
IV.3 $N=3$ Type 5 state (Fig. 1 (c))
@@
We can obtain such a state with an EPM bigraph Fig. 1 (c), whose corresponding optical circuit is presented in Fig. 5. Considering each dashed box as a dot, we can see that the wires are connected following the structure of the $N=3$ Type 5 entangled state sculpting bigraph. See Appendix C for the explicit calculation. $P_{suc}=5/(3^{2}2^{7})$ with feed-forward. To our knowledge, no heralded scheme for the $N=3$ Type 5 state has been proposed before.
@@
V Discussions
By establishing a set of translation rules from sculpting bigraphs into linear optical circuits, we have proposed a systematic method to design heralded circuits for the multipartite entanglment of bosons by combining the results in Ref. [16] with translation rules in Fig. 2. Although our work primarily addresses optical circuits employing polarization encoding, it is worth noting that conceptually equivalent circuits in dual-rail encoding can be directly designed (Apendix D).

### v3

@@ Abstract
Non-destructive heralded entanglement with photons is a valuable resource for quantum information processing. However, they generally entail ancillary particles and modes that amplify the circuit intricacy. To address this challenge, a recent work (npj Quantum Information 10, 67 (2024)) introduced a graph approach for creating multipartite entanglements with boson subtractions. Nonetheless, it remains an essential intermediate step toward practical heralded schemes: the proposition of heralded subtraction operators in bosonic linear quantum networks. This research establishes comprehensive translation rules from subtraction operators to linear optical operators, which provides a seamless path to design heralded schemes with single photons. Our method begets enhanced or previously unreported schemes for the $N$-partite GHZ state with $2N$ photons, $N$-partite W state with $12N+$ 1 photons and superposition of $N=3$ GHZ and W states with 9 photons. Our streamlined approach can straightforwardly design heralded schemes for multipartite entangled states by assembling the operators according to the guidence of sculpting bigraphs, hence significantly simplifies the quantum circuit design process.
1 Introduction
@@
In contrast, heralded schemes [11, 12, 13, 14, 15] adopts a different strategy by incorporating ancillary photons and modes as “heralds" to signal the successful generation of the anticipated target states. This allows for the discrimination of experimental runs that produce the desired target states without directly measuring those states. As a result, the entanglement generated through heralded schemes can be harnessed as a useful resource for quantum computations. However, the heralded schemes generally demand additional particles and modes for the resource of heralding, making them more intricate to design than postselected ones.
To overcome this difficulty of heralded schemes, a graph approach to heralded entanglement generation is proposed in Ref. [16], providing a systematic strategy to design the schemes by boson subtractions. Ref. [16] imposed correspondence relations between elements of bipartite graphs (bigraphs) and many-particle systems including boson annihilation operators, which is an extension of the graph picture introduced in Ref. [8]. To distinguish from other graphical approaches in quantum information and computation [17, 18, 19, 5], we call our approach the graph picture of linear quantum networks, or linear quantum graph (LQG) picture.
Based on the LQG picture, Ref. [16] provided various schemes to generate the multipartite entanglement of bosons with the sculpting protocol [20, 21], which generates an $N$-partite entangled state by applying tailored $N+K$ single-boson subtraction operators (which is dubbed the sculpting operator) with $K$ ancillary bosons. In the LQG picture, all the sculpting operators are represented as bigraphs (which are dubbed sculpting bigraphs). By leveraging the advantageous mathematical properties of the graphs, Ref. [16] has found essential entanglement generation schemes including qubit $N$-partite GHZ and W states, $N=3$ Type 5 entangled states (the superposition of $N=3$ GHZ and W states [22]), and qudit $N$-partite GHZ states.
@@
This work is organanized as follows: Section 2 explains the concept of sculpting protocol and its linear quantum graph (LQG) picture. Section 3 provides the translation rules from sculpting bigraphs to linear optical elements with the introduction of heralded subtraction operators (subtractors) in linear optics, by which we can construct heralded linear optical circuits that create GHZ, W, and the superposition of $N=3$ GHZ and W states in Section 4. Section 5 summarized the physical implications of our research and further applications.
2 Sculpting protocol and its linear quantum graph (LQG) picture
@@ 2.1 Sculpting protocol
    $\quad(k^{(l)}_{j,s}\in\mathbb{C}\leavevmode\nobreak\ \textrm{and}\leavevmode\nobreak\ \sum_{j,s}|k^{(l)}_{j,s}|^{2}=1).$ (3)
@@
2.2 Directed bipartite graph representation of bosonic systems in LQG picture
In Ref. [16], an undirected bigraph mapping was introduced (see Appendix A for a brief summary of mapping), which was useful for finding several sculpting schemes in the work. One can consult Appendix A of Ref. [8] for a simple glossary to the graph theory. While the undirected bigraph representation includes all the crucial information on the sculpting operators that can generate entanglement, we need a more comprehensive $directed$ bigraph representation to embrace the initial state and the sculpting operator in the LQG picture, which was first suggested in Appendix A of Ref. [16] and will be more thoroughly explained here. We construct all our schemes based on the directed bigraph representation in this paper as it has a more direct and intuitive relation with heralded linear optical circuits. In the directed bigraph picture, the complete information on the sculpting protocol from the preparation to the final state is encoded in a graph, which we call a sculpting directed bigraph.
@@ 2.2 Directed bipartite graph representation of bosonic systems in LQG picture
The first identity of Eq. (2.2) is expressed with directed bigraphs as
@@
[Uncaptioned image]= 0.[Uncaptioned image] $\end{gathered}=\leavevmode\nobreak\ 0.$ (10)
@@
2.3 EPM directed bigraphs for generating entanglement
@@
A crucial property of an EPM bigraph is that the final state generated by the corresponding sculpting operator is fully determined by the perfect matchings (PMs) of the bigraph (a set of disjoint edges that cover all the vertices in a given bigraph, see Appendix A of Ref. [8]), which we can show using identities (8) and (10[16]. We can directly check that all the sculpting bigraphs for generating genuine qubit entanglement in Ref. [16] (see (44)) correspond to EPM bigraphs.
@@
Comparing the EPM directed bigraphs in Fig. 1 with the undirected EPM bigraphs in (44), we can easily see that the right hand side of the directed ones in Fig. 1 (a), (b), and (c) without direction are identical to the three graphs in (44) of Appendix A. This is by the fact that the information on the sculpting operators is on the right hand side of the directed graphs, which is solely represented in the undirected graphs in (44). This shows that the essential part in a EPM directed bigraph is the right hand side of the graph from the circles as it represents the sculpting operator that generates entanglement.
Remark 3.—
It is also worth mentioning the fundamental conceptual difference between the tensor network notation [28, 18] and our LQG picture. The tensor network notation is a graphical mapping of distinguishable qudits (states and effects) and operations (processes) among them that can correspond to some type of nonlocality. On the other hand, the physical elements that the LQG picture represents are discrete spatial modes and identical particles (bosons or fermions) that has some dynamical relations to them, as shown in Tables 1. These elements cannot be considered qubit or qudit until they are encoded within the context of quantum information. For example, the initial state denoted as Eq. (6) cannot be expressed as a multi-qubit state. We can state that the LQG picture describes crude physical elements that can be combined to play the role of qubits.
3 Translation rules from sculpting bigraph elements to linear optical elements
Ref. [16] exploited the special property of EPM bigraphs (Property 2 of Section IV in Ref. [16]) to find schemes for genuinely entangled states, which are enumerated in Fig. 1. To design counterparts in linear optical circuits to the EPM bigraphs that generate the same entagled states, we need to know how each element of the graphs is translated to linear optical operators. A prerequisite for the translation rules from graphs to circuits is to provide heralded linear optical operators (which we name subtractors) that play the role of subtraction operators, because they are the fundamental components of EPM sculpting bigraphs.
@@ 2. $K$ circles to which edges are attached as
3.2 Subtraction operators in linear optics with heralding
@@ 2. Half-wave plates with HWP: $\{H,V\}\leftrightarrow\{D,A\}$ and HWP: $\{H,V\}\leftrightarrow\{A,D\}$
3. $n$-partite multiports that Fourier-transform the spatial mode states of photons (a $2$-partite port corresponds to a beam splitter (BS))
@@
$\{|0\rangle,|1\rangle\}\cong\{|D\rangle,|A\rangle\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \{|+\rangle,|-\rangle\}\cong\{|H\rangle,|V\rangle\}.$ (24)
@@
    ${\hat{a}}^{\dagger}_{H}{\hat{a}}^{\dagger}_{V}\to{\hat{a}}^{\dagger}_{D}{\hat{a}}^{\dagger}_{A}=\frac{1}{2}({\hat{a}}^{\dagger 2}_{H}-{\hat{a}}^{\dagger 2}_{V})$ (25)
@@
    $\frac{1}{2}({\hat{a}}^{\dagger 2}_{1,H}-{\hat{a}}^{\dagger 2}_{2,V})\to{\hat{a}}^{\dagger}_{1,H}.$ (27)
    Here a photon number resolving detector is required to execute the postselection of the one-photon state. We can see that the measurement of one photon at the detector 12 corresponds to the subtraction of one photon.
@@
where the incoming wires are attached to $11$ and $22$ by the internal states of the annihilation operators. The beam splitter (denoted as BS) is always balanced in our work and hence considered the 2-partite port. In the above subtractor, the two open wires go back to the spatial modes from which those wires came and be merged by PBSs as in (23) and (29).
@@
With the subtractors as building blocks, we can enumerate all the elements that construct EPM directed bigraphs and present translation rules of them to linear optical elements. All the translation rules from EPM sculpting bigraph elements to linear optical elements that we have explained are enumerated in Fig. 2. Therefore, an EPM bigraph that generates a multipartite entangled state serves as a blueprint for designing a linear optical circuit that generates the same entangled state, as will be clarified in the next section.
@@
4 Linear optical circuits
@@
Following the translation rules, we simply obtain a linear optical circuit for the GHZ state as Fig. 3. Comparing the circuit in Fig. 3 with the GHZ sculpting bigraph (Fig. 1 (a)), we can see that the photon paths are connected following the structure of the bigraph. The explicit calculation for showing that the above circuit generates the GHZ state is given in Appendix B. The success probability $P_{suc}$ is $\frac{1}{2^{2N-1}}$ with feed-forward [^2] The feed-forward scheme involves measurement result dependent single qubit operations, which are not challenging quantum operations in themselves. However, the scheme requires fast data processing to apply these operations in real-time based on measurement outcomes [30]. .
We can compare our scheme with other heralded GHZ-state-generating schemes [23, 24, 31] using various criteria. Before we get into the discussion, it is worth noting that all the schemes mentioned here require the same number of single photons for the same number of parties ($2N$ photons for the $N$-partite GHZ state). First, our scheme can be directly generalized to any $N$-partite case unlike Ref. [24] that suggests schemes for $N=2$ and 3 based on numerical optimizations or Ref. [23] for the $N=3$ GHZ entanglement generation using fusion gates (see Refs. [32, 33] for the experimental results). Second, contrary to schemes in Refs. [24, 31] that depend on the dual-rail encoding structure, our scheme can be installed with any kind of qubit encoding (polarization, dual-rail, etc.) based on the sculpting bigraph. See Appendix E for the transformation from polarization encoding circuits to dual-rail encoding circuits. Finally, we can also compare the optical component complexity [^3] It is worth noting that a direct comparison of the complexity or number of elements among different schemes is subtle in this case. The biggest reason is that each scheme requires a different type of devices, e.g., polarizing beam splitters (PBS) or beam splitters (BS) of various transmissivities. In addition, Refs. [23,24] is just about the N=3 GHZ state scheme, while ours presents a scheme for an arbitrary N.. Our scheme requires $3N$ PBSs and $2N$ detectors, which is simpler than the scheme in Ref. [31] that requires $4N$ BSs and $2N$ detectors. Additionally comparing with the schemes in Refs. [23, 24] for $N=3$ case, Ref. [23] requires the simplest circuit elements (5 PBSs and 6 detectors) and Ref. [24] the most complicated (12 BSs of various transitivities and 6 detectors).
4.2 $N$-partite W state (Fig. 1, (b))
@@
A linear optical scheme for the W generating EPM bigraph (Fig. 1, (b)) from the translation rules is as in Fig. 4. Considering each dashed box (spatially overlapped subtractors) as a dot, we can see that the wires are connected following the structure of the W sculpting bigraph. Appendix C provides an explicit calculation for showing that the above circuit generates the W state with $P_{suc}=\frac{1}{2^{2N}}$ with feed-forward.
To our knowledge, no linear optical heralded W-state scheme for arbitrary $N$ party with single-photon sources has been proposed. Since our scheme requires $12N+$ 1 single-photons, it is much more feasible than those schemes which exploit fusion gates and small size of W states as resources for generating $N$-partite W states [25, 26].
4.3 $N=3$ Type 5 state (Fig. 1 (c))
@@
We can obtain such a state with an EPM bigraph Fig. 1 (c), whose corresponding optical circuit is presented in Fig. 5. Considering each dashed box as a dot, we can see that the wires are connected following the structure of the $N=3$ Type 5 entangled state sculpting bigraph. See Appendix D for the explicit calculation. $P_{suc}=5/(3^{2}2^{7})$ with feed-forward. To our knowledge, no linear optical heralded scheme for the $N=3$ Type 5 state has been proposed before.
@@
5 Discussions
By establishing a set of translation rules from sculpting bigraphs into linear optical circuits, we have proposed a systematic method to design heralded circuits for the multipartite entanglment of bosons by combining the results in Ref. [16] with translation rules in Fig. 2. Although our work primarily addresses optical circuits employing polarization encoding, it is worth noting that conceptually equivalent circuits in dual-rail encoding can be directly designed (Appendix E).
Our GHZ scheme provides a general framework for an arbitrary number of parties and is compatible with any form of qubit encoding. It also does not have any trusted third-party for heralding the target state, which is advantageous in quantum communication. See Ref. [34] for a more detailed explanation on this aspect of our GHZ scheme. Ref. [34] also provides an analysis on the effect of channel losses on our scheme. For the cases of W and $N=3$ Type 5 state, no linear optical heralded scheme has been proposed to our knowledge. Given the latest experimental demonstrations [32, 33] of generating the $N=3$ GHZ state based on the scheme in Ref. [23], we believe that our proposed scheme is also realizable in the near future. Alternatively, one can also use online quantum computing platforms such as Quandela [32] that offer the linear operations involving 6 ideal photons with the integrated photonics by adopting our circuits into dual-rail encoding circuits.